# SEEK: Semantic Reasoning for Object Goal Navigation in Real World Inspection Tasks

### v1

@@ I Introduction
Performing object-goal navigation in unseen environments presents three main challenges. The first challenge is the large search space for the robot to consider when deciding where to go and inspect. The target object may be located in one room among many rooms in a building, and thus, uninformed brute-force search can be significantly inefficient. The second challenge is the robot’s inherent inability to use experience and common sense knowledge. Unlike humans, who draw upon prior knowledge and contextual understanding to refine their search strategies, the current state-of-the-practice object search methods are confined to sensory information onboard the robot. The third challenge is detecting the object under perceptual uncertainty due to the limited sensing range, inter-object occlusion, and detection and localization errors of the perception model on the robot. For example, planning for object search without probabilistic reasoning of the object detection accuracy could lead to inefficient search policy when there are false positives of the object detection [25, 63]. Decision making without probabilistic reasoning of prior semantic knowledge and perceptual uncertainty can lead to a myopic greedy policy or an overly conservative inefficient policy. These challenges underscore a fundamental barrier in object-goal navigation, highlighting the need for a new framework that mimics human’s reasoning capabilities.
@@
To address these challenges, we develop a framework that uses prior environment information and semantic knowledge in a probabilistic manner. While standard object-goal navigation problems only assume access to robot sensory readings [3], our approach can improve robot search efficiency by leveraging both i) prior spatial configuration and ii) relational semantic knowledge that are commonly accessible to humans performing inspection tasks. The environment spatial configuration, such as room layouts or blueprints, and Building Information Modeling (BIM) [54], provides a structured context for the robot to search the environment hierarchically. Additionally, relational semantic knowledge, such as typical locations of objects within the environment, can guide the object search. For instance, knowing that fire extinguishers are often found near entrances or kitchens, we can direct the robot to search those regions to locate the object. The advent of Large Language Models (LLMs) and language embeddings  [9, 18] allows us to access this semantic knowledge without human assistance. By leveraging both spatial and semantic information, we can formulate a probabilistic optimization problem for more efficient and robust object-goal navigation under uncertainty. We solve for a non-myopic policy which prioritizes the region with a higher probability of containing the target object as well as takes into account the chance of not finding the target object in the selected region, its associated costs, and the best follow-up policy in a recursive fashion. This capability can set a new standard for object-goal navigation in real-world inspections.
@@ II Related Works
Planning for autonomous inspection: Our approach is related to planning methods for autonomous robot exploration and inspection in real-world environments. Coverage planning is a well-studied method to explore and map environments efficiently [8, 39, 10, 13, 4]. In addition to map the whole environment, inspection tasks often require the robot to closely examine specific target objects of interest in some part of the environment [29]. Current state-of-the-practice approaches to inspecting specific targets usually rely on predefining routes and observation points for the robot or placing identifiable tags, such as Apriltags [56] or QR codes [6], making the process labor-intensive for humans. Recent approaches to addressing this problem involve training an object detector and incorporating the detection model into the planning process [16, 26, 19, 25]. For instance, the SWAP planner [19] integrates volumetric exploration with semantics coverage and inspection behavior, while the SB2G [25] initially employs geometric coverage behavior and transitions to active semantic search when the robot detects the target object. However, these methods are often inefficient in locating the object, particularly in an environment with a large search space, because these methods make the robot cover all unexplored areas in the environment until it detects the target object. In our work, we use a semantic-guided global planner instead of geometric coverage when the robot does not detect the target object. We use relational semantic knowledge to guide the robot to search promising regions in the environment.
@@
Semantic knowledge representation: Semantic knowledge equips robots with richer contextual information. Robots gather semantic information from their surroundings using techniques like object detection [42] and semantic segmentation [27]. This information can be represented in various forms, such as volumetric voxel maps [26], neural radiance fields (nerfs)[34], or 3D dynamic scene graphs (DSGs)[44]. In our work, we use a DSG to model the environment. The DSG’s structure aligns well with our hierarchical planning method, which utilizes multiple spatial abstraction layers within the DSG. The DSG can be constructed using blueprints because of the similarity in their structures. Moreover, the advancements in LLMs and vision-language models [15, 40] have led to the new paradigm of scene representations enriched with comprehensive semantic concepts and an extensive open-set object vocabulary [32, 46, 30]. Embedding human knowledge about semantic concepts and their interrelations enhances reasoning capabilities in planning. While open-vocabulary representations can augment our planning approach, we opt for a lightweight prediction model. This decision is driven by the need to deploy the model on robots with limited computational power and restricted internet access, especially in industrial inspection sites. In this work, we propose an object relationship representation that we call a Relational Semantic Network (RSN). The RSN distills knowledge from LLMs about the likelihood of encountering objects in various regions and refines its predictions based on the robot’s observations in the specific environment. Given the prior knowledge of the environment, the RSN is trained to predict objects relevant in environment-specific applications and can infer semantically related objects outside of the training data distilled from LLMs.
@@
Semantic-guided object-goal navigation is a prominent research area within embodied AI [3], where advancements are driven by the availability of realistic datasets [11, 61], simulators [49, 17], and competitive challenges [62]. Leading approaches train end-to-end reinforcement learning policies (RL) [37, 55] or combine RL with classical planning methods [12, 41]. Recently, zero-shot approaches have gained popularity through the use of pre-trained language or vision-language models (VLMs), as gathering robotics navigation data is resource-intensive [36, 23]. While LLMs can be directly queried to make decisions [48, 60], their planning and reasoning capabilities still raise uncertainties [53]. Recent state-of-the-art approaches use LLMs and VLMs with a frontier selection approach [14, 47, 64, 63]. Shah et al. employ LLMs as a guiding heuristic to bias exploration in the direction where the target object may be located [47]. Similarly, Chen et al. present frontier-based exploration with semantic utilities computed using language embeddings [14]. Yokoyama et al. query VLMs to generate a language-grounded value map by and choose the most promising frontier [63]. More related to our work, Zhou et al. reason over the semantic relationship between objects and rooms and formulates frontier selection with probabilistic soft logic [64]. Our work parallels these methods by incorporating common sense knowledge into the planning, but it is distinct in its approach by focusing on a room-based global planning policy, a slightly different problem setup that is particularly suited for urban inspection scenarios that have access to the spatial configuration of the environment.
@@
Statement of contribution: Our work makes three key contributions in the context of the current literature. Unlike state-of-the-art approaches that primarily rely on selecting the best semantic frontier for navigation [14, 47], our method introduces a probabilistic planning framework as an MDP, enabling a non-myopic strategy that plans actions over a longer horizon. This allows for consideration of future consequences and potential information gains, aiming to compute a globally optimal policy rather than just focusing on immediate short-term outcomes. Furthermore, we extend the use of semantic knowledge and the environment information mapped by the robot by introducing a new object-goal navigation problem setup, where the robot leverages the hierarchical structure of environments provided by human blueprints. Our framework also aims to bridge the gap in applying semantic-based object-goal navigation in practical applications. To the best of our knowledge, our work is the first approach that leverages both prior spatial configurations from blueprints and relational semantic knowledge for probabilistic planning in real-world object-goal navigation. The practicality of this approach on inspection tasks is demonstrated on hardware, highlighting its relevance and applicability in real-world settings.
@@
Figure 2: SEEK system architecture. Our method receives an inspection query and builds a Dynamic Scene Graph (DSG) from a given blueprint. The Relational Semantic Network (RSN) predicts the probability of finding the target object in every room. The semantic-guided global planner uses this prediction to compute an optimal global search policy. Finally, the local finite state controller executes the global policy and switches to an active semantic search controller when the robot sees the object.
@@ III-A Problem Statement
Object-goal navigation: The robot’s objective is to search and navigate to a target object $y_{G}\in Y$ for object inspection. The target object is described as the object semantic class $y^{l}$. The robot successfully navigates to $y_{G}$ when the robot position $x_{k}^{p}$ is close to the target object $\|x_{k}^{p}-y_{G}^{p}\|<\epsilon$ and triggers a finish signal $u^{f}$. The robot must efficiently search for the object by minimizing the distance travelled to the target object.
@@ III-B Semantic Knowledge Representation
We use two sources of prior semantic knowledge to guide the robot to locate the target object. The first source of prior knowledge comes from the spatial configuration of the environment. This information can come from blueprints and Building Information Models (BIM) commonly available in industrial inspection tasks. The spatial configuration provides the names of regions or rooms in the environment that are semantically meaningful to locate the target object. The second source of prior knowledge is the semantic relationship between target objects and room names. The semantic relationship informs which rooms the target object is usually located in. We maintain spatial configuration in a Dynamic Scene Graph (DSG) and semantic knowledge relationship in a Relational Semantic Network (RSN).
@@ III-C Planning with Relational Semantic Knowledge
where $\mathcal{X}_{i}$ and $\mathcal{X}_{j}$ denotes the current state and the next state, respectively.
@@ IV Semantic-guided Object-Goal Navigation
Having outlined our approach for solving object-goal navigation with relational semantic knowledge, we discuss how we build the DSG and RSN and compute the object-goal navigation policy. Our approach consists of an offline initialization of the DSG and RSN and an online object-goal planning. Before deploying the robot for the task, we generate the DSG from blueprints (subsection IV-A). We then train the RSN to predict object occurrences to guide the robot in searching for objects that can be found in the environment (subsection IV-B). During the deployment, when the robot receives a request to inspect a target object $y_{G}$, the global planner computes the global policy that minimizes the expected travel distance to search for $y_{G}$ through the rooms or regions in the environment (subsection IV-C). Finally, the local finite state controller selects a control policy for the robot based on the estimated belief of the target object (subsection IV-D). As the robot observes the environment, it recomputes its global and local policy and saves the DSG and RSN for future deployments. Fig. 2 summarizes the architecture of SEEK.
@@ IV-A Dynamic Scene Graph (DSG) Generation
We generate the DSG of the environment using the environment’s blueprint. Blueprints provide geometric and semantic information about the room/region nodes $\mathcal{V}_{2}$. For the location nodes $\mathcal{V}_{1}$, we sample the nodes uniformly across all the open space based on the blueprint to ensure full room coverage. We compute the Euclidean distance between $\mathcal{V}_{1}$ and store the information on the edges.
When the robot is being deployed in the environment, the robot updates the DSG. First, the robot self-localizes using its LIDAR-based map to the blueprint [43]. The robot navigates in the environment and updates the traversability risks [22] between the location nodes. The connectivity can change across deployments due to the evolving object configuration in the environment over time that is not described in the blueprints. During the deployment and in our experiments, we do not consider changes in the connectivity between rooms and buildings. When the robot observes and localizes an object, a new object node is created on the first layer and connected to the nearest location node. The updated DSG is used for path planning and stored for future deployments.
@@ IV-C Semantic-Guided Global Planner
Given the DSG and the RSN, we have all the information to solve the semantic-guided global planning problem through the MDP formulation. The state space $\mathbb{X}$ and action space $\mathbb{U}$ are extracted from the set of room nodes $\mathcal{V}_{2}$ in the DSG. To specify the transition probabilities $P$, we first set the transition probabilities to reach the goal state $P(\mathcal{X}_{k+1}=\mathcal{X}^{G}\mid\mathcal{X}_{k},\mathcal{U}_{k})$ by querying the RSN. The transition between rooms is deterministic since the DSG $\mathcal{G}$ is a connected graph. Then we can set the transition probabilities to reach a new room after taking $\mathbb{U}^{move}$ action and staying at the same room after taking $\mathbb{U}^{\text{search}}$ action as $1-P(\mathcal{X}_{k+1}=\mathcal{X}^{G}\mid\mathcal{X}_{k},\mathcal{U}_{k})$. To specify the cost function $C$, we compute the shortest travel distance to visit rooms using A search [21] on the DSG and compute the total distance to visit location nodes on every room. Given that the room connectivity is assumed not to change during the run, the cost function $C$ can be computed offline and stored as a look-up table for planning. The cost function can be recomputed during the run if we consider a robot deployment with changing room connectivity. The cost function can be recomputed between the global planning computation.
@@ IV-D Local Finite State Controller
We use three different controllers that execute the global policy $\pi^{g}$, search and navigate to the object. We use the Semantic Belief Behavior Graph (SB2G) framework to implement and define the controller switching [25].
Global search controller $\rho_{\text{nav}}$ computes a robot policy according to the global policy $\pi^{g}(x_{k})$. If the global action is to move to a new room $\mathbb{U}^{\text{move}}$, the controller computes the shortest path to the room on the graph and passes the path to a risk-aware MPC controller that computes a robot trajectory [22]. If the global action is to search the current room, the controller plans a coverage path that explores the obstacle free-area in the room [8] and passes the path to the MPC controller.
@@ V-A Global Planner Evaluation
We demonstrate the performance of our approach in the Habitat simulator using the Matterport3D dataset [11]. We are not training our RSN with any data from the dataset. The Matterport3D dataset provides a hierarchical scene configuration that we use to build the DSG. The robot is tasked to navigate to the closest object instance [3]. We sample 5 scenes and 10 different objects from the Matterport3D dataset. The simulations were performed on a laptop with an Intel i9-11950H CPU.
@@
TABLE I: Global planning results. We report the mean and standard deviation of the Success weighted by inverse path length (SPL). Results are separated into two categories: Fixed Objects and Movable Objects, to demonstrate the effectiveness of different methods on different type of objects.
@@
Table I presents the performance of various methods at locating fixed objects (e.g., bed, sink) and movable objects (e.g., clothes, towel). Our method outperforms other methods on both object categories. Our approach is significantly better in locating fixed objects, which can be attributed to their distinct association with specific room types. In contrast, searching for movable objects is relatively more challenging as they can be found in various rooms.
@@
Figure 4: Robot paths comparison of object-goal navigation with SEEK and SB2G in the Gazebo simulator. Our global planner guides the robot to the kitchen where a fire extinguisher is usually located.
Figure 5: The left plot compares how fast the robot search and navigate to the target object from the same starting location using SEEK and SB2G. The shades on the line plot represent the standard deviation across three runs with the same starting location. The right box plot reports the SPL metric across 7 runs with different starting location.
@@
Figure 7: Experiment results on hardware. The left figure visualizes the global planning policy executed by the robot. We marks the room with high probability of finding the target object based on the RSN prediction. The middle figure plots the path taken by the robot. The right figure shows third person view of the robot on different phase of the inspection.

### v2

# This paper has been published in the proceedings of Robotics: Science and Systems, 2024. Please cite this paper as:

@@ I Introduction
Performing object-goal navigation in unseen environments presents three main challenges. The first challenge is the large search space for the robot to consider when deciding where to go and inspect. The target object may be located in one room among many rooms in a building, and thus, uninformed brute-force search can be significantly inefficient. The second challenge is the robot’s inherent inability to use experience and common sense knowledge. Unlike humans, who draw upon prior knowledge and contextual understanding to refine their search strategies, the current state-of-the-practice object search methods are confined to sensory information onboard the robot. The third challenge is detecting the object under perceptual uncertainty due to the limited sensing range, inter-object occlusion, and detection and localization errors of the perception model on the robot. For example, planning for object search without probabilistic reasoning of the object detection accuracy could lead to inefficient search policy when there are false positives of the object detection [25, 63]. Decision making without probabilistic reasoning of prior semantic knowledge and perceptual uncertainty can lead to a myopic greedy policy or an overly conservative inefficient policy. These challenges underscore a fundamental barrier in object-goal navigation, highlighting the need for a new framework that mimics human reasoning capabilities.
@@
To address these challenges, we develop a framework that uses prior environment information and semantic knowledge in a probabilistic manner. While standard object-goal navigation problems only assume access to robot sensory readings [3], our approach can improve robot search efficiency by leveraging both i) prior spatial configuration and ii) relational semantic knowledge that are commonly accessible to humans performing inspection tasks. The environment spatial configuration, such as room layouts, floor plans, blueprints, and Building Information Modeling (BIM) [54], provides a structured context for the robot to search the environment hierarchically. Additionally, relational semantic knowledge, such as typical locations of objects within the environment, can guide the object search. For instance, knowing that fire extinguishers are often found near entrances or kitchens, we can direct the robot to search those regions to locate the object. The advent of Large Language Models (LLMs) and language embeddings  [9, 18] allows us to access this semantic knowledge without human assistance. By leveraging both spatial and semantic information, we can formulate a probabilistic optimization problem for more efficient and robust object-goal navigation under uncertainty. We solve for a non-myopic policy that prioritizes the region with a higher probability of containing the target object as well as takes into account the chance of not finding the target object in the selected region, its associated costs, and the best follow-up policy in a recursive fashion. This capability can set a new standard for object-goal navigation in real-world inspections.
@@ II Related Works
Planning for autonomous inspection: Our approach is related to planning methods for autonomous robot exploration and inspection in real-world environments. Coverage planning is a well-studied method to explore and map environments efficiently [8, 39, 10, 13, 4]. In addition to map the whole environment, inspection tasks often require the robot to closely examine specific target objects of interest in some part of the environment [29]. Current state-of-the-practice approaches to inspecting specific targets usually rely on predefining routes and observation points for the robot or placing identifiable tags, such as Apriltags [56] or QR codes [6], making the process labor-intensive for humans. Recent approaches to addressing this problem involve training an object detector and incorporating the detection model into the planning process [16, 26, 19, 25]. For instance, the SWAP planner [19] integrates volumetric exploration with semantics coverage and inspection behavior, while the SB2G [25] initially employs geometric coverage behavior and transitions to active semantic search when the robot detects the target object. However, these methods are often inefficient in locating the object, particularly in an environment with a large search space, because these methods make the robot cover all unexplored areas in the environment until it detects the target object. In our work, we use a semantic-guided global planner instead of geometric coverage when the robot does not detect the target object. We use relational semantic knowledge to guide the robot in searching promising regions in the environment.
@@
Semantic knowledge representation: Semantic knowledge equips robots with richer contextual information. Robots gather semantic information from their surroundings using techniques like object detection [42] and semantic segmentation [27]. This information can be represented in various forms, such as volumetric voxel maps [26], neural radiance fields (NeRF)[34], or 3D dynamic scene graphs (DSGs)[44]. In our work, we use a DSG to model the environment. The DSG’s structure aligns well with our hierarchical planning method, which utilizes multiple spatial abstraction layers within the DSG. The DSG can be constructed using floor plans because of the similarity in their structures. Moreover, the advancements in LLMs and vision-language models [15, 40] have led to the new paradigm of scene representations enriched with comprehensive semantic concepts and an extensive open-set object vocabulary [32, 46, 30]. Embedding human knowledge about semantic concepts and their interrelations enhances reasoning capabilities in planning. While open-vocabulary representations can augment our planning approach, we opt for a lightweight prediction model. This decision is driven by the need to deploy the model on robots with limited computational power and restricted internet access, especially in industrial inspection sites. In this work, we propose an object relationship representation that we call a Relational Semantic Network (RSN). The RSN distills knowledge from LLMs about the likelihood of encountering objects in various regions and refines its predictions based on the robot’s observations in the specific environment. Given the prior knowledge of the environment, the RSN is trained to predict objects relevant to environment-specific applications and can infer semantically related objects outside of the training data distilled from LLMs.
@@
Semantic-guided object-goal navigation is a prominent research area within embodied AI [3], where advancements are driven by the availability of realistic datasets [11, 61], simulators [49, 17], and competitive challenges [62]. Leading approaches train end-to-end reinforcement learning policies (RL) [37, 55] or combine RL with classical planning methods [12, 41]. Recently, zero-shot approaches have gained popularity through the use of pre-trained language or vision-language models (VLMs), as gathering robotics navigation data is resource-intensive [36, 23]. While LLMs can be directly queried to make decisions [48, 60], their planning and reasoning capabilities still raise uncertainties [53]. Recent state-of-the-art approaches use LLMs and VLMs with a frontier selection approach [14, 47, 64, 63]. Shah et al. employ LLMs as a guiding heuristic to bias exploration in the direction where the target object may be located [47]. Similarly, Chen et al. present frontier-based exploration with semantic utilities computed using language embeddings [14]. Yokoyama et al. query VLMs to generate a language-grounded value map by and choose the most promising frontier [63]. More related to our work, Zhou et al. reason over the semantic relationship between objects and rooms and formulate frontier selection with probabilistic soft logic [64]. Our work parallels these methods by incorporating common sense knowledge into the planning, but it is distinct in its approach by focusing on a room-based global planning policy, a slightly different problem setup that is particularly suited for urban inspection scenarios that have access to the spatial configuration of the environment.
@@
Statement of contribution: Our work makes three key contributions in the context of the current literature. Unlike state-of-the-art approaches that primarily rely on selecting the best semantic frontier for navigation [14, 47], our method introduces a probabilistic planning framework as an MDP, enabling a non-myopic strategy that plans actions over a longer horizon. This allows for consideration of future consequences and potential information gains, aiming to compute a globally optimal policy rather than just focusing on immediate short-term outcomes. Furthermore, we extend the use of semantic knowledge and the environment information mapped by the robot by introducing a new object-goal navigation problem setup, where the robot leverages the hierarchical structure of environments provided by floor plans. Our framework also aims to bridge the gap in applying semantic-based object-goal navigation in practical applications. To the best of our knowledge, our work is the first approach that leverages both prior spatial configurations from floor plans and relational semantic knowledge for probabilistic planning in real-world object-goal navigation. The practicality of this approach on inspection tasks is demonstrated on hardware, highlighting its relevance and applicability in real-world settings.
@@
Figure 2: SEEK system architecture. Our method receives an inspection query and builds a Dynamic Scene Graph (DSG) from a given floor plan. The Relational Semantic Network (RSN) predicts the probability of finding the target object in every room. The semantic-guided global planner uses this prediction to compute an optimal global search policy. Finally, the local finite state controller executes the global policy and switches to an active semantic search controller when the robot sees the object.
@@ III-A Problem Statement
Object-goal navigation: The robot’s objective is to search and navigate to a target object $y_{G}\in Y$ for object inspection. The target object is described as the object semantic class $y^{l}$. The robot successfully navigates to $y_{G}$ when the robot position $x_{k}^{p}$ is close to the target object $\|x_{k}^{p}-y_{G}^{p}\|<\epsilon$ and triggers a finish signal $u^{f}$. The robot must efficiently search for the object by minimizing the distance traveled to the target object.
@@ III-B Semantic Knowledge Representation
We use two sources of prior semantic knowledge to guide the robot in locating the target object. The first source of prior knowledge comes from the spatial configuration of the environment. This information can come from floor plans and Building Information Models (BIM) commonly available in industrial inspection tasks. The spatial configuration provides the names of regions or rooms in the environment that are semantically meaningful in locating the target object. The second source of prior knowledge is the semantic relationship between target objects and room names. The semantic relationship informs which rooms the target object is usually located in. We maintain spatial configuration in a Dynamic Scene Graph (DSG) and semantic knowledge relationship in a Relational Semantic Network (RSN).
@@ III-C Planning with Relational Semantic Knowledge
where $\mathcal{X}_{i}$ and $\mathcal{X}_{j}$ denote the current state and the next state, respectively.
@@ IV Semantic-guided Object-Goal Navigation
Having outlined our approach for solving object-goal navigation with relational semantic knowledge, we discuss how we build the DSG and RSN and compute the object-goal navigation policy. Our approach consists of an offline initialization of the DSG and RSN and an online object-goal planning. Before deploying the robot for the task, we generate the DSG from floor plans (subsection IV-A). We then train the RSN to predict object occurrences to guide the robot in searching for objects that can be found in the environment (subsection IV-B). During the deployment, when the robot receives a request to inspect a target object $y_{G}$, the global planner computes the global policy that minimizes the expected travel distance to search for $y_{G}$ through the rooms or regions in the environment (subsection IV-C). Finally, the local finite state controller selects a control policy for the robot based on the estimated belief of the target object (subsection IV-D). As the robot observes the environment, it recomputes its global and local policy and saves the DSG and RSN for future deployments. Fig. 2 summarizes the architecture of SEEK.
@@ IV-A Dynamic Scene Graph (DSG) Generation
We generate the DSG of the environment using the environment’s floor plan. Floor plans provide geometric and semantic information about the room/region nodes $\mathcal{V}_{2}$. For the location nodes $\mathcal{V}_{1}$, we sample the nodes uniformly across all the open space based on the floor plan to ensure full room coverage. We compute the Euclidean distance between $\mathcal{V}_{1}$ and store the information on the edges.
When the robot is being deployed in the environment, the robot updates the DSG. First, the robot self-localizes using its LIDAR-based map to the floor plan [43]. The robot navigates in the environment and updates the traversability risks [22] between the location nodes. The connectivity can change across deployments due to the evolving object configuration in the environment over time, which is not described in the floor plans. During the deployment and in our experiments, we do not consider changes in the connectivity between rooms and buildings. When the robot observes and localizes an object, a new object node is created on the first layer and connected to the nearest location node. The updated DSG is used for path planning and stored for future deployments.
@@ IV-C Semantic-Guided Global Planner
Given the DSG and the RSN, we have all the information to solve the semantic-guided global planning problem through the MDP formulation. The state space $\mathbb{X}$ and action space $\mathbb{U}$ are extracted from the set of room nodes $\mathcal{V}_{2}$ in the DSG. To specify the transition probabilities $P$, we first set the transition probabilities to reach the goal state $P(\mathcal{X}_{k+1}=\mathcal{X}^{G}\mid\mathcal{X}_{k},\mathcal{U}_{k})$ by querying the RSN. The transition between rooms is deterministic since the DSG $\mathcal{G}$ is a connected graph. Then we can set the transition probabilities to reach a new room after taking $\mathbb{U}^{move}$ action and staying in the same room after taking $\mathbb{U}^{\text{search}}$ action as $1-P(\mathcal{X}_{k+1}=\mathcal{X}^{G}\mid\mathcal{X}_{k},\mathcal{U}_{k})$. To specify the cost function $C$, we compute the shortest travel distance to visit rooms using A search [21] on the DSG and compute the total distance to visit location nodes in every room. Given that the room connectivity is assumed not to change during the run, the cost function $C$ can be computed offline and stored as a look-up table for planning. The cost function can be recomputed during the run if we consider a robot deployment with changing room connectivity. The cost function can be recomputed between the global planning computation.
@@ IV-D Local Finite State Controller
We use three different controllers that execute the global policy $\pi^{g}$, search, and navigate to the object. We use the Semantic Belief Behavior Graph (SB2G) framework to implement and define the controller switching [25].
Global search controller $\rho_{\text{nav}}$ computes a robot policy according to the global policy $\pi^{g}(x_{k})$. If the global action is to move to a new room $\mathbb{U}^{\text{move}}$, the controller computes the shortest path to the room on the graph and passes the path to a risk-aware MPC controller that computes a robot trajectory [22]. If the global action is to search the current room, the controller plans a coverage path that explores the obstacle-free area in the room [8] and passes the path to the MPC controller.
@@ V-A Global Planner Evaluation
We demonstrate the performance of our approach in the Habitat simulator using the Matterport3D dataset [11]. We are not training our RSN with any data from the dataset. The Matterport3D dataset provides a hierarchical scene configuration that we use to build the DSG. The robot is tasked to navigate to the closest object instance [3]. We sample five scenes and ten different objects from the Matterport3D dataset. The simulations were performed on a laptop with an Intel i9-11950H CPU.
@@
TABLE I: Global planning results. We report the mean and standard deviation of the Success weighted by inverse path length (SPL). Results are separated into two categories: Fixed Objects and Movable Objects, to demonstrate the effectiveness of different methods on different types of objects.
@@
Table I presents the performance of various methods at locating fixed objects (e.g., bed, sink) and movable objects (e.g., clothes, towel). Our method outperforms other methods on both object categories. Our approach is significantly better at locating fixed objects, which can be attributed to their distinct association with specific room types. In contrast, searching for movable objects is relatively more challenging as they can be found in various rooms.
@@
Figure 4: Robot paths comparison of object-goal navigation with SEEK and SB2G in the Gazebo simulator. Our global planner guides the robot to the kitchen, where a fire extinguisher is usually located.
Figure 5: The left plot compares how fast the robot searches and navigates to the target object from the same starting location using SEEK and SB2G. The shades on the line plot represent the standard deviation across three runs with the same starting locations. The right box plot reports the SPL metric across seven runs with different starting location.
@@
Figure 7: Experiment results on hardware. The left figure visualizes the global planning policy executed by the robot. We mark the rooms with a high probability of finding the target object based on the RSN prediction. The middle figure plots the path taken by the robot. The right figure shows a third-person view of the robot on different phases of the inspection.